Cash Budget Daybook Inc. collects 30% of its sales on account in the month of the sale and 70% in the month following the salProduction Budget Daybook Inc. projected sales of 400,000 personal journals for 20Y6. The estimated January 1, 20Y6, inventorDirect Materials Purchases Budget Daybook Inc. budgeted production of 403,500 personal journals in 2016. Paper is required toDirect Labor Cost Budget Daybook Inc. budgeted production of 403,500 personal journals in 20Y6. Each journal requires assemblCost of Goods Sold Budget Daybook Inc. budgeted production of 403,500 personal journals in 2016. Paper is required to produceCost of direct materials placed in production $ Total work in process during period
Answer #1

A 1 2 Cash Budget 3 Budgeted Cash receipts from sales on account for October 4. 5 September Sales 6 October Sales 105000 1160

Cash Budget
Budgeted Cash receipts from sales on account for October
September Sales               1,05,000
October Sales               1,16,000
30% Same month 34800
70% following month 73500
Total Cash receipts budgeted 108300
Production Budget Units
Sales               4,00,000
Add: Closing Stock desired                  23,500
Less: Opening Inventory                 -20,000
Budgeted Production              4,03,500
Particulars Square yards
Opening Inventory 40400
Add: Purchases (Balancing figure) 2419500
Less: Closing inventory desired -38900
Production (403500 journals * 6 yards for each journal) =403500*6
Direct Materials Purchases Budget =2419500*0.4
=$ 967,800

Ε Time required in minutes =403500*8 Time required in hours =+E4/60 Direct Labor Cost =53800*13

Time required in minutes 3228000
Time required in hours 53800
Direct Labor Cost        6,99,400